OpenStudy (anonymous):

log base 7 (x+4) + log base 7 (x-2) =1

OpenStudy (anonymous):

exact same idea as before

OpenStudy (anonymous):

i know it just didn't work out for me this time

OpenStudy (anonymous):

1) combine into one log \[\log_7((x+4)(x-2))=1\]

OpenStudy (anonymous):

2)rewrite in equivalent exponential form: \[(x+4)(x-2)=7\]

OpenStudy (anonymous):

3) solve the resulting quadratic equation

OpenStudy (anonymous):

discard any answers that would make the input of your logarithm negative

OpenStudy (anonymous):

...

OpenStudy (anonymous):

log base 7 (x+4) + log base 7 (x-2) =1 log base 7 (x+4)(x-2)=1 x^2+2x-8=7 x^2+2x-15=0 x^2+5x-3x-15=0 (x+5)(x-3)=0 x=-5,x=3

OpenStudy (anonymous):

nvm i am having trouble concentrating! i get the questions sorry for the trouble of answering them.. maybe i should pay attention to what im writing

OpenStudy (anonymous):

yeas and don't forget to discard -5 because it does not work!

OpenStudy (anonymous):

Yeah I made a point about antilogarithms in your previous post. The always have to be greater than zero or the same way satellite put it you have to discard any solutions that make the input of your log zero. One way to do it is to set up the anti logs like this: x + 4 > 0 , and x - 2 >0, When you solve this x > 2 so when you get a solution like x = -5 it doesn't satisfy this condition and therefore that solution is not true.

OpenStudy (anonymous):

want some medals?.. \[\log_{2}(2m+4)-\log_{2}(m-1)=3\]